One description of the potential energy of a diatomic molecule is given by the Lennard-Jones potential. U= A/r^12 - B/r^6 where A and B are constants and r is the separation distance between the atoms. For the H2 molecule, take A=0.124 x 10^12eV m12 and B=1.488 x 10^-60eV m6. Find (a) the separation distance r0 at which the energy of the molecule is a minimum and (b) the energy E required to break up the H2 molecule.
1
Expert's answer
2020-10-27T08:23:00-0400

Given expression for Potential is,

"U=\dfrac{A}{r^{12}}-\dfrac{B}{r^6}~~~~~~-(1)"


(a) When the energy is minimum ,it means the system is in equlibrium

Differentiate equation 1 with respect to r

"\to \dfrac{dU}{dr}=\dfrac{-12A}{r^{13}}+\dfrac{6B}{r^7}"


For energy minimum "\dfrac{du}{dr}=0"


"\therefore \dfrac{12A}{r^{13}}=\dfrac{6B}{r^7}"



"\to r^6=\dfrac{2A}{B}"


"\to r^6=\dfrac{2\times0.124\times 10^{12}}{1.488\times 10^{-60}}"


"\to r^6=\dfrac{0.248\times10^{72}}{1.488\times 10^{-60}}"


"\to r^6=0.16\times 10^{72}"


so r="(0.16\times 10^{72})^{\frac{1}{6}}" m


(b) Energy required to brake "H_2" molecule is

"\to U=\frac{0.124\times 10^{12}}{0.16\times10^{72}}-\frac{1.488\times 10^{-60}}{0.16\times10^{72}}"


=0.775-9.3"\times 10^{-132}" Joule